a train takes 3 hours and 10 minutes to reach its destination .approximately how many hours does the trip take

Answer by nyc_function(2741)

In travel problems, time is measured in terms of hours not minutes.

So, 10 minutes = 10/60 = 1/6 of an hour.

Then 3 + 1/6 = 3_1/6 hours.

We know that 3_1/6 is a mixed number that can be changed to an improper fraction.

So, 3_1/6 = 19/6.

Dividing 19 by 6, we get 3.166, which can be rounded to the nearest tenths place to become 3.2 hours.